Two forces act on an object, producing a resultant force of 356.2 pounds. One of the forces if 200 pounds and makes an angle of 47 degrees with the resultant force. Approximate the magnitude of the otehr force to the nearest pound.
Answer by jojo14344(1512)

We should remember the total resultant force= 356.2lbs (Ft)
F1 + F2 = 356.2 lbs ------------- eqn 1
Since, a tangential force makes an angle 47deg of force 200lbs, we need to get the force on the "x" (Fx) and "y" (Fy) axis to get the total forcce for that particular impact. To show in eqn, Fx + Fy = F1 ------- eqn 2
For Fx, we need to use trigo function cosine since it is adjacent to the force 200lbs at an angle 47 deg= (Fx) cos47deg.
Also for Fy, we need to use sine since it is opposite the force 200lbs at an angle 47deg= (Fy)sine47deg. Going back eqn 2,
(Fx)cos47 + (Fy)sin47 = F1
(200)cos47 + (200)sin47 = F1
282.67lbs = F1
So for the other force F2, go back eqn 1,
282.67 + F2 = 356.2
F2=73.53 = 74lbs (the other you're looking for)
Thank you,
Jojo